OpenStudy (anonymous):

f(x)={(3,5), (2,4), (1,7)} what does 1/f(x)=

OpenStudy (anonymous):

Each given point is of the form \(\left(x,~f(x)\right)\). You want points of the form \(\left(x,~\dfrac{1}{f(x)}\right)\). To do this, you keep the x-coordinates the same, but take the reciprocal of the y-coordinates. So if \(f(x)=\{(3,5),~(2,4),~(1,7)\}\), then \(\dfrac{1}{f(x)}=\left\{\left(3,?\right),~\left(2,?\right),~\left(1,?\right)\right\}\).